p=1/2mv^2 solve for m

Answer by ptaylor(2198)

p=1/2mv^2 solve for m
p=(1/2)mv^2 multiply both sides by 2
2p=mv^2(2/2)
2p=mv^2 divide both sides by v^2
2p/(v^2)=m


Hope this helps-------ptaylor